Cooper Ltd. has 2 operating divisions: domestic sales and international sales. They also have 2 support divisions: accounting support and human resources support. For the past year, Cooper’s cost records show the following information:

Support Divisions

Operating Divisions

Account’g Support

Human Resources Support

Domestic Sales

Inter-national Sales

Total

Budgeted costs incurred before any interdivision cost allocations

$500,000

$600,000

$8,400,000

$7,500,000

$ 17,000,000

Support work supplied by Accounting (based on # of employees)

20%

40%

40%

100%

Support work supplied by Human Resources (based on # of staffing actions)

10%

60%

30%

100%

Required:

  1. Allocate the 2 support divisions’ costs to the operating divisions using the Direct Method.
  2. Allocate the 2 support divisions’ costs to the operating divisions using the Reciprocal Method.
  3. Allocate the 2 support divisions’ costs to the operating divisions using the Step Down Method – Allocate Accounting Support first.
  4. What method do you recommend and why?

Cooper Ltd. would like to share the services of a consultant with another company, Hofstadter Inc. The consultant will work a total of 1,000 hours for both companies: 700 for Cooper and 300 hours for Hofstadter Inc. The total cost of the contract will be $368,000. If Cooper were to contract directly with the consultant for 700 hours, the cost would have been $350,000. If Hofstadter were to contract directly with the consultant, the contract would have cost them $ 150,000. How should the companies divide up the cost of the consultant?
Answer #1
Part (a)
Support Divisions Operating Divisions
Account’g Support Human Resources Support Domestic Sales Inter-national Sales Total
Departmental cost before allocation $500,000 $600,000 $8,400,000 $7,500,000 $17,000,000
Allocations
Accounts Depatment cost
- equally given to each operartion depatments
'- i.e. 1/2 , 1/2
($500,000) $250,000 $250,000
HR Departemnt cost
'- 6/9, 3/9
($600,000) $400,000 $200,000

*Department Accounts’s cost has been allocated on the basis of no of employees:

40%+ 40% = 80%.

i.e ½ , ½

Allocated to department Domestic: $500,000 × (1/2) = $250,000
Allocated to department International : $500,000 × (1/2) = $250,000


**Department HR cost has been allocated on the basis of STAFFING ACTION:

60 + 30

    90

Allocated to department Domestic: $600,000 × (6/9) = $400,000
Allocated to department International : $600,000 × (3/9) = $200,000

Answere Part (b)

in image formate Ans (c)

Sr.no. Support Divisions Operating Divisions
1 Account’g Support Human Resources Support Domestic Sales Inter-national Sales
2 Departmental cost before allocation $500,000 $600,000 $8,400,000 $7,500,000
3 Allocations
4 Accounts Depatment cost
- equally given to each operartion depatments
'- i.e. 1/2 , 1/2
($500,000) $100,000 $200,000 $200,000
5 Total $0.00 $700,000.00 $8,600,000.00 $7,700,000.00
6 HR Departemnt cost
'- 6/9, 3/9
$0.00 ($700,000) $466,666.67 $233,333.33
7 Total = 5+6 $0.00 $9,066,666.67 $7,933,333.33

Answere (d)

I Will always recommend Reciprocal method because  the direct method completely ignores the interdepartmental services and step method gives them only a partial recognition because it allocates costs forward – never backward.

To overcome for the same, Reciprocal is the best.